How long would it take a machine to do 10,000 joules of work if the power rating of the machine is 100 watts?


1
Expert's answer
2021-03-14T19:14:14-0400
"P=\dfrac{W}{t},""t=\dfrac{W}{P}=\dfrac{10000\ J}{100\ W}=100\ s."